If 0.106 moles of zinc reacts with excess lead(IV) sulfate, how many grams of zinc sulfate would be produced?
1
Expert's answer
2016-03-07T08:44:14-0500
2Zn + Pb(SO4)2 = 2ZnSO4 + Pb
N(ZnSO4)=N(Zn)= 0.106 moles.
m(ZnSO4)=0.106* 161.45=17.11g